From nch@roe.ac.uk Fri Dec 7 14:53:49 2007 Date: Fri, 7 Dec 2007 14:15:12 +0000 (GMT) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Johann Bryant , Mike Read , Nigel Hambly , Nicholas Cross , Bob Mann , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Brian Walshe , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Mark Holliman , Peredur Williams , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 7/12/07 Minutes of WFAU VDFS Science Archive meeting: 7th December 2007 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, RSC, MAR, PMW, JB, ETWS, LGR, MSH, NJC Apologies: BCW, JPE, AL, RGM, JMS DONM: 10am, Friday 14th December 2007 NB: in the VISTA HUT Actions discharged this week: ----------------------------- ACTION: ALL to note changes and enhancements in dr3plus_release.ht Discharged ACTION: ETWS to import cirdr into the SVN as a separate folder under VDFSarchives but outside trunk. Discharged Actions partly discharged but continuing: ----------------------------------------- The following from last time partly done but continue: o) Add in a default row for every detector appearing in every detection table (for schema consistency when querying merged sources and individual detections) - ACTION: RSC & NCH Continues; will get sorted (honestly!) as part of the general ingest DB schema revamp to facilitate rapid generalised photometric/astrometric recalibration after DR3. ACTION: JB to complete operations "brain dump" onto the TWiki. Actions carried forward from 30/11/07 meeting: ---------------------------------------------- ACTION: JB & ETWS to restore SegueDR6 on thutmose at their convenience (i.e. no hurry given the present circumstances) Continues; team agreed to postpone until after DR3 now; also awaits instance of SQL 2005 on the batch catalogue server which in turn awaits full SSA tests against user interface JDBC applications... we're getting there... ACTION: JB to set dfg_1 shrinking in the WSA this weekend. Continues; will try this weekend end as DXS DR3 preparations are on hold (see below) Specific points and new actions: -------------------------------- Project management: PMW noted that it is that time again when we should fill in progress for November; the team filled in the various stats during the meeting. NCH noted that ESO Public Surveys Panel has sent around time allocation notices to some of the PIs, along with a doc specifying requirements for delivery of data products to the ESO "Science Archive" Facility. VDMT are reviewing these requirements at the suggestion of JPE to make sure the implications for VDFS are understood and planned for. The requirements will be factored into plans in the New Year, although NCH noted that given the current financial climate there is no guarantee that development funds will be forthcoming in the current Rolling Grant renewals. WFCAM & VISTA updates: No updates this week. Comments and issues arising from CASU fortnightly minutes: No new minutes this week. Networking: ETWS and JB noted that Sept 07B is now ok-to-copy at CASU (NCH also noted that MJI had confirmed this yesterday). Transfers will be instigated after this weekend since the power is going off to all servers in C1 for infrastructure works. WSA Operations: DR3 (phase 1) is out the door, thank the gods; backups of ingest DB and release DB have been secured by JB. GPS release preparations are going well (so far...) Hardware NCH noted that a single channel on one RAID card failed mid-week, just at the right time for DR3; Eclipse fixed it very promptly though. As regards the new catalogue SQL2005 server hatshepsut, NCH noted that the SSA load is progressing well but that we now need to fully integrate it into the archive private network. ACTION: MSH to bring hatshepsut onto private network, and to set up a login for MAR ACTION: ETWS to put FreeTDS DNS entries in to enable ODBC/JDBC connectivity to hatshepsut. MSH and JB noted the power downtime in server room C1 this weekend; plans are in place for the rearrangement of the archive hardware. Service will start to be interupted from somewhere between 4pm and 5pm this evening. Software: RSC reported: "I've been holding CU19's hand to make sure it could correctly produce a DR3 database without the GPS. Likewise making sure that CU7 is correctly source merging the GPS in stages. In my free time I've been attempting to bring the DEV1 branch back up to date, which has lied stagnant whilst DR3 software enhancements took priority. I need to close out the DEV1 branch, where I'll implement a proper fix to the index problem, before we can start work on the new schema changes to improve recalibration performance. ETWS reported: "Fixed minor problems with the DR3 webpage parsing and created webpages and metdata for Astrogrid. Started GPS DR3 release curation runs of CU7. Created svn entry for cirdr code. Reparsed constraints for sql tables to include missing cascading deletes." NJC reported: "This week I have been testing and putting together plots for the UKIDSS meeting in Munich; also made slight modifications to code to produce input files for CU8 to correct for the extension number error in CU4." Survey Data Release: DR3 (phase 1, i.e. without the GPS) was released on the great unwashed yesterday. GPS source merging is continuing apace. Non-survey Data Release: MAR reported three non-survey registrations in this week Astrogrid deployment: MSH and ETWS have completed registry metadocs for 2XMM and UKIDSS DR3; MSH noted that multiple source tables are now included for UKIDSS cone-searches. There was some debate over the incomplete and out-of-date UCD information currently provided in the WFAU schemas, but action on this was not felt to be urgent since it does not seem to be stopping any VO user applications from running (as far as we know). Miscellaneous: NCH noted that SJW has asked that UKIDSS Workshop presentations be made to stimulate discussion with the users, and NCH suggested we have a run-through of what we've got next week ACTION: NCH, ETWS, LGR and NJC to review their UKIDSS Workshop contributions next Wednesday at 2pm. Plate Library final removals will take place on Thurs 13th December. Early doors tonight (to celebrate DR3 and to give JB a good send-off) at 6pm in the Old Bell (within staggering distance to JB's gaff). Late-taste advanced rewarning for Fri 14th, Cannon's Gait from 8pm(ish) onwards for an Xmas p***-up featuring The Blackford Jazz Experience (good grief...)